# Machine Learning-Driven Insights into Excitonic Effects in 2D Materials

### v1

@@ 1 Introduction
Two-dimensional materials (2DMs) have garnered substantial interest due to their exceptional optical and electronic properties, positioning them as promising candidates for next-generation optoelectronic technologies. A defining feature of these materials is their pronounced excitonic effects, which are significantly more pronounced in 2D systems compared to bulk materials [1]. This enhanced excitonic behavior arises from reduced dielectric screening and increased Coulomb interactions in monolayers [2, 3]. Consequently, understanding these excitonic properties is crucial for optimizing the performance of devices based on such materials.
@@
First-principles calculations of exciton binding energy (EBE) in 2DMs involve computationally expensive and laborious methods, such as the GW approximation and the Bethe-Salpeter Equation (BSE) [4]. The GW approximation surpasses mean-field, independent-particle DFT by accounting for many-body electron-electron interactions, offering a more accurate understanding of electronic properties, including excitation energies, band gaps, and optical characteristics. This method involves iteratively solving for the Green’s function ($\mathbf{G}$), the screened Coulomb interaction ($\mathbf{W}$), and the self-energy ($\mathbf{\Sigma}$) until self-consistency is achieved, yielding an improved description of the electronic structure [5]. Following the GW calculation, the BSE method is employed to compute excitonic effects by incorporating electron-hole interactions, providing precise estimates of EBE through electron-hole correlation [3]. However, these methods are computationally demanding, making them less practical for large-scale material screening.
To tackle this challenge, we propose a machine learning-assisted approach to efficiently predict EBE. This approach takes advantage of data from first-principles calculations, such as band gaps obtained from PBE and HSE06 as presented in C2DB [6, 7, 8].
@@
Machine learning (ML) approach requires a fine balance between accuracy and computational efficiency in predicting complex quantities from simpler, readily available parameters [9]. By linking these simpler features to exciton binding energy estimates—which would otherwise require advanced computational methods—our approach accelerates the materials discovery process. This study highlights the potential of ML to streamline the identification of 2D materials with large excitonic effects, enabling faster screening for optoelectronic applications.
@@ 2.1 Dataset
In the context of machine learning in materials science, material databases are essential for the success of predictive models. Using a large amount of high-quality data is mandatory to achieve a robust and accurate predictions. In this study, we used C2DB database to train and evaluate our machine learning models, which comprises band gap data for around 4001 monolayers. These materials include transition metal dichalcogenide, transition metal oxides and other technologically important materials like hexagonal boron nitride [10], MgI2, MgBr2 [11] etc.
@@ 2.2 Features selection
The Pearson correlation coefficient quantifies the linear relationship between variables, providing insight into how one property influences another in monolayer materials. In this study, it serves as a critical tool to evaluate the dependencies between different features, A high Pearson coefficient indicates a strong linear association, such as between PBE and HSE06 band gaps, suggesting that trends in simpler PBE calculations can predict more computationally intensive HSE06 results. Similarly, correlations between G0W0 band gaps and other properties reflect how electronic interactions evolve between computational methods. Understanding these relationships helps optimize feature selection for predicting EBE.
The wrapper method evaluates the predictive power of features by assessing their impact on model performance. Unlike statistical filter methods, it iteratively trains a model with different feature subsets, capturing both linear and nonlinear dependencies. In this study, the wrapper method identified band gap values for PBE and HSE06 as the most significant predictors of EBE.
@@ 3.1 Predicting Quasi-particle energies
The Random Forest regression model demonstrates strong performance in predicting G0W0 band gaps, which correspond to quasiparticle (QP) energies. The model’s effectiveness is illustrated in Figure 1. The distribution of prediction errors for the G0W0 band gaps, using RF algorithm, demonstrates a strong alignment with actual values from the
@@
C2DB database, indicating the model’s high predictive accuracy and reliability. The histogram of errors has a clear peak around zero, as shown in Figure 2, indicating that most of the predictions are very close to the actual band gap values. This means that the RF model is accurate and effectively captures the relationships in the data.
@@ 3.2 Exciton binding Energy
The exciton binding energy is the energy needed to separate an exciton—a bound state of an electron and a hole—into free, independent charge carriers. EBE can be estimated using the band gaps obtained from HSE06 and G0W0 methods [18, 19] as:
$EBE=E_{g}\ (G_{0}W_{0})-E_{g}\ (HSE06)$ (1)
The EBE prediction trained using the RF algorithm is presented in Fig. 3, demonstrating that the this model effectively captures the EBE with high accuracy.
@@
The distribution of prediction errors for EBE, shown for the RF model in Fig. 4, exhibits a sharp peak at zero, highlighting minimal deviation between predicted and actual values and underscoring the model’s precision.
Figure 4: Distribution of prediction errors for EBE for RF, with a peak at zero, indicating high accuracy with minimal deviation between predicted and actual values.
The radar plots in Figure 5 illustrate the performance of five ML algorithms i.e. Gradient Boosting, Random Forest, Support Vector Regression, Kernel Ridge Regression, and Neural Networks—across training and testing sets, evaluated by mean absolute error (MAE), root mean square error (RMSE), and R2 metrics.
These metrics collectively assess each model’s prediction accuracy and capacity to generalize to new data set. In both training and testing sets, the RF model demonstrates superior performance, achieving a training MAE of 0.011 and RMSE of 0.025. Its testing metrics are also impressive, with MAE of 0.021 and RMSE of 0.043, indicating strong generalization with a minimal increase in error on unseen data. R2 scores of 0.994 signifies that it captures nearly all variance in the exciton binding energy as shown in Fig. 5.
@@
Gradient Boosting model follows closely, with an R2 of 0.981 as well in other statistical parameters as illustrated in Fig. 5. The Neural Network, KRR, and SVR algorithms relatively under-perform having R2 of around 0.95. Random Forest is particularly effective in modeling exciton binding energy due to its ensemble nature, which combines multiple decision trees to reduce over fitting and improve generalization. It captures complex, nonlinear relationships between the features and the target variable with high accuracy, as indicated by its $R^{2}$ of 0.994. RF also handles feature interactions well and is robust to noise, making it suitable for datasets with moderate complexity, as demonstrated by its consistently low MAE and RMSE on both training and test sets.
Lin et al. [20] used features such as the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the C2DB database to predict exciton binding energies in 2D materials. Their model achieved an $R^{2}$ value of 0.802 and a mean absolute error (MAE) of 0.214 eV with a gradient boosting regression model. Our RF-based model improves on these results by incorporating exciton-related features, achieving better $R^{2}$ and MAE values.
The derived linear regression equation for EBE demonstrates its dependence on the PBE and HSE06 band gaps.
$EBE=0.48\ E_{PBE}+0.01\ E_{HSE06}+0.28$ (2)
The equation indicates that PBE contributes significantly to predicting the EBE, highlighting its dominant role. In contrast, the HSE06 band gap has a minimal influence, as reflected by its small coefficient of 0.01. The constant term 0.28 accounts for the base offset in EBE values. This model provides a simple, interpretable relationship between the EBE and the electronic properties i.e. PBE and HSE06 band gap methods, enabling a straightforward estimation of EBE using cheaply available features.
@@ 4 Conclusion
In this study, we demonstrated a machine learning-assisted approach for predicting exciton binding energies in two-dimensional materials, using band gap data from simple DFT calculations. By training and evaluating multiple ML algorithms, we found that the RF model provided the most reliable predictions, effectively bridging the gap between computational efficiency and accuracy. Our model, trained on PBE and HSE06 band gap data from the C2DB, offers a rapid and cost-effective alternative to traditional GW and BSE methods, enabling faster screening and discovery of materials with significant excitonic effects. Additionally, we implemented a Bayesian optimization framework, which further streamlined the identification of top EBE monolayers by efficiently guiding the search for promising candidates. This integration of BO with machine learning underscores the potential for data-driven approaches to revolutionize materials discovery processes.
5 Author contributions
A. Javed conceived the idea, developed initial concepts and performed the calculations and analysis. A. Sajid supervised the project and helped editing the manuscript.

### v2

@@ 1 Introduction
Two-dimensional materials (2DMs) have garnered substantial interest due to their exceptional optical and electronic properties, positioning them as promising candidates for next-generation optoelectronic technologies. A defining feature of these materials is their pronounced excitonic effects, which are significantly more pronounced in 2D systems compared to bulk materials [1, 2]. This enhanced excitonic behavior arises from reduced dielectric screening and increased Coulomb interactions in monolayers [3, 4]. Consequently, understanding these excitonic properties is crucial for optimizing the performance of devices based on such materials.
The exciton binding energy (EBE) is a fundamental property of excitons, defined as the energy required to dissociate a bound electron-hole pair (exciton) into free charge carriers. In first-principles calculations, EBE determination typically requires computationally intensive methods such as the GW approximation and then Bethe-Salpeter equation (BSE) [5]. The EBE is then calculated as:
$EBE=E_{g}^{electronic}-E_{g}^{optical}$ (1)
@@
Here $E_{g}^{electronic}$ is the GoWo-corrected quasi-particle band gap, $E_{g}^{optical}$ is energy of the lowest-energy bright exciton from BSE [6]. The GW approximation surpasses mean-field, independent-particle DFT by accounting for many-body electron-electron interactions, offering a more accurate understanding of electronic properties, including excitation energies, band gaps, and optical characteristics. This method involves iteratively solving for the Green’s function ($\mathbf{G}$), the screened Coulomb interaction ($\mathbf{W}$), and the self-energy ($\mathbf{\Sigma}$) until self-consistency is achieved, yielding an improved description of the electronic structure [7]. Following the GW calculation, the BSE method is employed to compute excitonic effects by incorporating electron-hole interactions, providing precise estimates of EBE through electron-hole correlation [4]. However, these methods are computationally demanding, making them less practical for large-scale material screening.
To address this challenge, we propose a machine learning-assisted approach for efficiently predicting EBE. This method leverages data such as band gaps obtained from PBE and structural parameters available in the C2DB database. By correlating these simpler features with exciton binding energy estimates, which would typically require advanced computational methods, this approach accelerates the materials discovery process.
While Liang et al. (2019) pioneered physics-inspired machine learning for band gap and EBE prediction based on Phillips ionicity theory, their reliance on small datasets and ad-hoc structural maps limits the generalizability of their approach. Additionally, their model only covers AmBn materials, whereas the C2DB has been significantly updated over the past five years [8]. In another study, Lin et al. (2023) used features such as the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the C2DB database to predict EBE in 2D materials [9]. Their model achieved a $R^{2}$ value of 0.80 and a mean absolute error (MAE) of 0.21 eV with a gradient boosting (GB) regression model. Other models have the potential to improve this accuracy, as we will demonstrate in this work. While machine learning has been employed to predict EBE, Bayesian Optimization remains unexplored in this context.
@@
Machine learning (ML) approaches require a fine balance between accuracy and computational efficiency when predicting complex quantities from simpler, readily available parameters [10]. This study highlights the potential of ML to streamline the identification of 2D materials with large excitonic effects, enabling faster screening for optoelectronic applications.
@@ 2.1 Dataset
In the context of machine learning in materials science, material databases are essential for the success of predictive models. Using a large amount of high-quality data is mandatory to achieve a robust and accurate predictions. In this study, we used C2DB database [11, 12] to train and evaluate our machine learning models, which comprises band gap data for transition metal dichalcogenides (e.g., MoS2) [13], transition metal oxides and other technologically important materials like hexagonal boron nitride [14], MgI2, MgBr2 [15] etc.
@@ 2.2 Features selection
Pearson correlation coefficient quantifies the linear relationship between variables, providing insight into how one property influences another in monolayer materials. In this study, it serves as a critical tool to evaluate the dependencies between different features [17, 18]. A high Pearson coefficient indicates a strong linear association, such as between PBE and HSE06 band gaps, suggesting that trends in simpler PBE calculations can predict more computationally intensive HSE06 results. Similarly, correlations between G0W0 band gaps and other properties reflect how electronic interactions evolve between computational methods. The wrapper method evaluates the predictive power of features by assessing their impact on model performance. Unlike statistical filter methods, it iteratively trains a model with different feature subsets, capturing both linear and nonlinear dependencies [19]. Understanding these relationships helps optimize feature selection for predicting EBE. Based on these methods, the features selected for this study include the layer thickness, layer group number, atomic number, valence electron count, spin, number of atoms and PBE band gap of the structure.
@@ 3.1 Predicting Quasi-particle energies
The Random Forest regression model demonstrates strong performance in predicting G0W0 band gaps, which correspond to quasiparticle (QP) energies. The model’s effectiveness is illustrated in Figure 1. The distribution of prediction errors for the G0W0 band gaps demonstrates a strong alignment with actual values from the C2DB database, indicating the model’s high predictive accuracy and reliability. The histogram of errors has a clear peak around zero, as shown in Figure 2, indicating that most of the predictions are very close to the actual band gap values, R2 equals to 0.98, with MAE and RMSE values of 0.20 and 0.33, respectively.
@@ 3.2 Exciton binding Energy
The exciton binding energy is the energy required to separate an exciton into free, independent charge carriers. The EBE prediction model, trained using the RF algorithm, is presented in Figure 3. This model effectively captures the EBE with high accuracy for the available data in the C2DB database, achieving an R2 value of 0.84.
@@
The radar plots in Figure 4 illustrate the performance of five ML algorithms i.e. Gradient Boosting, Random Forest, Support Vector Regression, Kernel Ridge Regression, and Neural Networks, across training and testing sets, evaluated by mean absolute error (MAE), root mean square error (RMSE), and R2 metrics.
These metrics collectively assess each model’s prediction accuracy and capacity to generalize to new data set. In both training and testing sets, the RF model demonstrates superior performance, achieving a training MAE of 0.070 and RMSE of 0.126. Its testing metrics are also impressive, with MAE of 0.081 and RMSE of 0.148, indicating strong generalization with a minimal increase in error on unseen data. R2 scores of 0.84 signifies that it captures most of the variance in the exciton binding energy as shown in Figure 4.
@@
Gradient Boosting model follows closely, with an R2 of 0.80 as well in other statistical parameters as illustrated in Fig. 4. The Neural Network, KRR, and SVR algorithms relatively under-perform having R2 of around 0.56. Random Forest is particularly effective in modeling exciton binding energy due to its ensemble nature, which combines multiple decision trees to reduce over fitting and improve generalization. It captures complex, nonlinear relationships between the features and the target variable with high accuracy. RF also handles feature interactions well and is robust to noise, making it suitable for datasets with moderate complexity, as demonstrated by its consistently low MAE and RMSE on both training and test sets.
While Lin et al. [9] employed HOMO and LUMO energies from the C2DB database as primary features for predicting EBE and achieving R2 of 0.80 using GB algorithm. Our Random Forest model demonstrates superior predictive performance through incorporation of additional physically meaningful descriptors and yielding improved accuracy with R2 equal to 0.84. This advancement highlights the importance of comprehensive feature engineering in excitonic property prediction.
The derived linear regression equation for EBE demonstrates its dependence on the dominating features.
$EBE=0.31\ E_{PBE}-0.15\ Z-0.06\ d+0.11\ n_{ve}+0.94$ (2)
The equation shows that the PBE band gap is a significant contributor to predicting the exciton binding energy, emphasizing its dominant role. The variable ‘Z’ represents the atomic number, ‘d’ stands for the thickness, and ‘$n_{ve}$’ indicates the valence electron count. This model provides a simple and interpretable relationship between the EBE and the most influential features.
@@ 4 Conclusion
In this study, we demonstrated a machine learning-assisted approach for predicting exciton binding energies in two-dimensional materials, using band gap data from simple DFT calculations. By training and evaluating multiple ML algorithms, we found that the RF model provided the most reliable predictions, effectively bridging the gap between computational efficiency and accuracy. Our model offers a rapid and cost-effective alternative to traditional GW and BSE methods, enabling faster screening and discovery of materials with significant excitonic effects. Additionally, we implemented a Bayesian optimization framework, which further streamlined the identification of top EBE monolayers by efficiently guiding the search for promising candidates. This integration of BO with machine learning underscores the potential for data-driven approaches to revolutionize materials discovery processes.
5 CRediT author statement
A. Javed: Conceptualization, Methodology, Software, Writing - Original Draft A. Sajid: Supervision, Validation, Writing - Review & Editing.